A solution containing 28% acid is to be mixed with a solution containing 40% acid to make a 300 L solution which contains 36% acid. How of each solution should be used to make the desired mixture?
Answer by checkley77(12844)

.40x+.28(300-x)=.36*300
.40x+84-.28x=108
.12x=108-84
.12x=24
x=24/.12
x=200 L of 40% solution is used.
300-200=100 L of 28% solution is used.
Proof:
.40*200+.28*100=.36*300
80+28=108
108=108